OpenStudy (anonymous):

Find the solution of 2x^2*y"+xy'-3y=0, y(1)=1, y'(1)=4.

OpenStudy (loser66):

Is that Euler's formula apply?

OpenStudy (anonymous):

I don't know, do I need to use the method that I used last time?

OpenStudy (loser66):

I'll be back, need eating something, give me time for eating,

OpenStudy (anonymous):

Okay.

OpenStudy (loser66):

read my note, part Euler's Equation, and try it first

OpenStudy (anonymous):

\(y=x^r\) checks out.

OpenStudy (anonymous):

So do you use that substitution again like last time to find the general solution?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wait a minute.

OpenStudy (anonymous):

So I got r=-1, 3/2. Now what?

OpenStudy (anonymous):

Now your general solution is \[y=C_1x^{3/2}+C_2x^{-1}\] Use the initial conditions to solve for the constants.

OpenStudy (anonymous):

So when I plug in y(1)=1, I got c1+c2=1.

OpenStudy (anonymous):

Right, and what about y'(1)?

OpenStudy (anonymous):

4=c1*3/2*x^(1/2)-c2*x^-2

OpenStudy (anonymous):

Replace x with 1... Anyway, you have the system \[\begin{cases}C_1+C_2=1\\\frac{3}{2}C_1-C_2=4\end{cases}\] Solve for the constants, replace them in the general solution, and you're done!

OpenStudy (anonymous):

Thank you, Siths, you made everything so easy.

OpenStudy (loser66):

@SithsAndGiggles we don't care about the coefficient of x? I mean 2 in this case.

OpenStudy (anonymous):

You're welcome. @Loser66, what do you mean? Did you find \(C_1=2\)?

OpenStudy (loser66):

nope, I talk about your first comment "y = x^r check out"

OpenStudy (anonymous):

Yes, that's the guess. At this point, you could include the arbitrary \(C\), but the algebra is cleaner without it. In the end, you have to remember that \(x^r\) is a particular solution and \(Cx^r\) is a general solution. Using the first is slightly more convenient.

OpenStudy (loser66):

on the previous post the coefficient of x is 1, you suggest let (x+c)^r because it form from the form (x+1) whose coefficient of x is 1, this case, it's 2

OpenStudy (anonymous):

Ah, I see what you mean. The coefficients in the ODE don't change the answer.

OpenStudy (loser66):

Got you. Thanks a lot. I can learn from other's post, hihihi... That's great!!